# Expansive Supervision for Neural Radiance Fields

### v2

@@ Abstract
Neural Radiance Field (NeRF) has achieved remarkable success in creating immersive media representations through its exceptional reconstruction capabilities. However, the computational demands of dense forward passes and volume rendering during training continue to challenge its real-world applications. In this paper, we introduce Expansive Supervision to reduce time and memory costs during NeRF training from the perspective of partial ray selection for supervision. Specifically, we observe that training errors exhibit a long-tail distribution correlated with image content. Based on this observation, our method selectively renders a small but crucial subset of pixels and expands their values to estimate errors across the entire area for each iteration. Compared to conventional supervision, our approach effectively bypasses redundant rendering processes, resulting in substantial reductions in both time and memory consumption. Experimental results demonstrate that integrating Expansive Supervision within existing state-of-the-art acceleration frameworks achieves 52% memory savings and 16% time savings while maintaining comparable visual quality.
@@ I Introduction
Despite the impressive performance of NeRF, its training speed remains a significant limitation. In the original NeRF architecture, rendering each pixel requires sampling $N$ points to compute the color $\mathbf{c}$ and density $\sigma$. For a scene with dimensions $(h,w)$, this process necessitates $h\cdot w\cdot N$ neural network forward passes, potentially exceeding $10^{8}$ computations for rendering a single 1080p resolution view. This computational burden substantially prolongs the training duration and impedes the practical generation of neural 3D media content. To address these computational challenges, various acceleration methods have been proposed, including decomposition schemes [2], distillation to light fields [3], and incorporation of explicit representations [4, 5]. Among these, the most effective approaches leverage explicit representations to cache view-independent features, trading memory consumption for reduced training time. These efficient explicit representations include voxel grids [6, 7], point cloud [8], hash tables [5], and low-rank tensors [4]. By incorporating such explicit structures, training time for a scene can be reduced significantly, from approximately 10 hours to 30 minutes.
@@ II-A Preliminary and Formulation
Figure 3: Pipeline of expansive supervision. The mechanism of expansive supervision is to exclusively render the crucial pixels, which consist of the pre-computed anchor area and sampled source areas, to estimate the loss. This estimation is accomplished through the expansive strategy described in Section II-B. Note that the expansive strategy is implemented on permuted pixels, i.e., random batch training in NeRF. Our expansive supervision method results in significant time and memory savings while maintaining negligible compromise in visual quality.
@@ III-A Implementation Details
Datasets & Metrics. We evaluate our method using both the object-centric Synthetic-NeRF [1] and real-world forward-facing LLFF [12] datasets. For quantitative assessment, we employ PSNR, SSIM [13], and LPIPS [14] metrics, where L(V) denotes the VGG [15] versions of LPIPS.

### v3

@@ Abstract
Neural Radiance Field (NeRF) has achieved remarkable success in creating immersive media representations through its exceptional reconstruction capabilities. However, the computational demands of dense forward passes and volume rendering during training continue to challenge its real-world applications. In this paper, we introduce Expansive Supervision to reduce time and memory costs during NeRF training from the perspective of partial ray selection for supervision. Specifically, we observe that training errors exhibit a long-tail distribution correlated with image content. Based on this observation, our method selectively renders a small but crucial subset of pixels and expands their values to estimate errors across the entire area for each iteration. Compared to conventional supervision, our approach effectively bypasses redundant rendering processes, resulting in substantial reductions in both time and memory consumption. Experimental results demonstrate that integrating Expansive Supervision within existing state-of-the-art acceleration frameworks achieves 52% memory savings and 16% time savings while maintaining comparable visual quality. Our code is available at this link.
@@ I Introduction
Despite the impressive performance of NeRF, its training speed remains a significant limitation. In the original NeRF architecture, rendering each pixel requires sampling $N$ points to compute the color $\mathbf{c}$ and density $\sigma$. For a scene with dimensions $(h,w)$, this process necessitates $h\cdot w\cdot N$ neural network forward passes, potentially exceeding $10^{8}$ computations for rendering a single 1080p resolution view. This computational burden substantially prolongs the training duration and impedes the practical generation of neural 3D media content. To address these computational challenges, various acceleration methods have been proposed, including decomposition schemes [2], distillation to light fields [3], and incorporation of explicit representations [4, 5]. Among these, the most effective approaches leverage explicit representations to cache view-independent features, trading memory consumption for reduced training time. These efficient explicit representations include voxel grids [6], hash tables [5], and low-rank tensors [4]. By incorporating such explicit structures, training time for a scene can be reduced significantly, from approximately 10 hours to 30 minutes.
@@ II-A Preliminary and Formulation
Figure 3: Pipeline of expansive supervision. The mechanism of expansive supervision is to exclusively render the crucial pixels, which consist of the pre-computed anchor area and sampled source areas, to estimate the loss. This estimation is accomplished through the expansive strategy described in Section II-B.
@@ III-A Implementation Details
Datasets & Metrics. We evaluate our method using both the object-centric Synthetic-NeRF [1] and real-world forward-facing LLFF [10] datasets. For quantitative assessment, we employ PSNR, SSIM, and LPIPS [11] metrics, where L(V) denotes the VGG [12] versions of LPIPS.
@@ IV Conclusion
Acknowledgment
National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China(Grant No. 92467204 and 62472249), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).